# LazyEviction: Lagged KV Eviction with Attention Pattern Observation for Efficient Long Reasoning

### v1

@@ 5 Experimental Results
In this section, we select DeepSeek-R1-Distill-Llama-8B  [^1] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B and DeepSeek-R1-Distill-Qwen-7B  [^2] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B as the base models for evaluating the efficiency of LazyEviction. The evaluations are carried on two widely used mathematical reasoning benchmarks: GSM8Kcobbe2021training and MATH-500hendrycks2021measuring . We compare LazyEviction with FullKV and three representative KV cache compression methods, including TOVA oren2024transformers , H2O zhang2023h2o , and RaaS hu2025efficient . More implementation details can be found in Appendix A.2.
@@ 5.1 Performance Comparison with Baselines
Table 1 presents a comprehensive performance comparison of our LazyEviction and baselines. The experimental results demonstrate that LazyEviction consistently outperforms other methods across two datasets and two models, achieving performance comparable to the baseline FullKV with minimal deviation. In DeepSeek-R1-Distill-Llama-8B, at a KV cache compression ratio of 50%, our method maintains performance comparable to FullKV, while other methods exhibit significant degradation. Notably, the RaaS algorithm shows an accuracy drop of nearly 50% on the GSM8K dataset. In DeepSeek-R1-Distill-Qwen-7B, at a KV cache compression ratio of 65%, our method achieves an accuracy reduction of less than 2.1% compared to FullKV, outperforming other three algorithms.
Table 1: Performance of our LazyEviction compared with three baselines on the GSM8K and MATH-500 dataset with DeepSeek-R1-Distill-Llama-8B and DeepSeek-R1-Distill-Qwen-7B. The best results among all methods are in bolded.
Base Model DeepSeek-R1-Distill-Llama-8B DeepSeek-R1-Distill-Qwen-7B
Dataset GSM8K MATH-500 GSM8K MATH-500
FullKV 74.68 62.0 91.74 86.4
KV Cache Comression Ratio = 50% KV Cache Comression Ratio = 65%
H2O zhang2023h2o 62.24 (-12.44) 56.0 (-6.0) 89.38 (-2.36) 81.2 (-5.2)
TOVA oren2024transformers 66.26 (-8.42) 56.2 (-5.8) 80.14 (-11.6) 79.0 (-7.4)
RaaS hu2025efficient 39.88 (-34.8) 58.2 (-3.8) 86.35 (-5.39) 83.0 (-3.4)
LazyEviction 72.70 (-1.98) 60.4 (-1.6) 90.07 (-1.67) 84.6 (-1.8)
Figure 5: Trade-off between accuracy and cache budget of LazyEviction and other baselines on (a) GSM8K and (b) MATH-500 datasets using DeepSeek-R1-Distill-Llama-8B Model, and (c) GSM8K and (d) MATH-500 datasets using DeepSeek-R1-Distill-Qwen-7B Model.
Trade-off between Accuracy and Cache Budget
We evaluating the performance of LazyEviction in optimizing the trade-off between KV cache budget and accuracy compared to baselines. As shown in Fig.  5, our method maintains higher accuracy across various budgets for both datasets and models. When the cache budget is large, other methods perform slightly worse than our approach. However, with a smaller cache budget, other methods experience substantial accuracy degradation. Notably, the RaaS algorithm experiences a pronounced accuracy drop at low cache budgets, as it prioritizes retaining prompt tokens, which limits the retention of other essential tokens and leads to poor performance. By implementing lagged KV eviction and leveraging MRI to track recurring tokens, our approach consistently achieves optimal performance, even under stringent budget constraints.
@@ 5.2 Memory Efficiency of LazyEviction
Figure 6: KV cache memory usage of different algorithms, including (a) the KV cache memory usage with varying output length (0-8k tokens) on DeepSeek-R1-Distill-Qwen-7B, and average KV cache memory usage on (b) DeepSeek-R1-Distill-Llama-8B and (c) DeepSeek-R1-Distill-Qwen-7B for GSM8K datasets under different cache budget.
We evaluate the KV cache memory usage of different algorithms. First, we evaluate the KV cache memory usage on the MATH-500 dataset using DeepSeek-R1-Distill-Qwen-7B, analyzing its variation with output token length. As shown in Fig. 6, the memory consumption of FULLKV increases linearly with the number of tokens. In contrast, the memory usage of TOVA, H20, and RaaS algorithms remains constant once the token count exceeds the budget. LazyEviction, with the observation window mechanism, exhibits minor fluctuations in KV cache memory usage after surpassing the budget, but Fig. 6 indicates that its memory growth is minimal compared to the other three algorithms. To further assess LazyEviction’s memory efficiency, we conduct experiments by calculating the average KV cache memory usage on the GSM8K dataset. Fig. 6 illustrates the performance of DeepSeek-R1-Distill-Llama-8B, demonstrating that our approach reduces cache memory to 55.85% while maintaining an accuracy drop of less than 1%. Similarly, Fig.  6 shows the performance of DeepSeek-R1-Distill-Qwen-7B, where our method achieves a 50.8% reduction in cache memory with an accuracy decrease of less than 5%.
@@ 5.3 Ablation Study
Table 2: Baseline algorithms with observation window mechanism, evaluating on the GSM8K dataset using DeepSeek-R1-Distill-Llama-8B with $B=192$ and $W=25$.
@@
Budget = 192
LazyEviction 73.62
H2O 66.26
+ window (W=25) 69.98 (+3.72)
TOVA 70.43
+ window (W=25) 71.34 (+0.91)
RaaS 54.36
+ window (W=25) 58.21 (+3.85)
@@
Since our method retains a slightly increased number of KV pairs within the observation window, which may lead to the perceived performance gains being predominantly attributed to these additional tokens. To eliminate the impact of the additional tokens within observation window, we explore the effects of other algorithms with the observation window mechanism on performance. As shown in Table 2, when H20, TOVA, and RaaS algorithms perform KV eviction every $W$ decoding steps, their performance improves. This improvement arises from the linear increase in retained KV pairs between consecutive decision decoding steps. However, their performance remains inferior to that of LazyEviction. LazyEviction employs MRI to track recurring tokens, effectively capturing their potential future importance, which leads to superior performance.
The Selection of Window Size $W$.
We investigate the impact of the window size $W$ on DeepSeek-R1-Distill-Llama-8B performance across two datasets in Table 3. As $W$ increases, the model’s accuracy generally improves. This is because a larger window size allows the LazyEviction mechanism’s MRI metric to detect more recurring tokens, resulting in better capturing a token’s potential future significance, thereby enhancing performance. However, when the window size becomes too large, LazyEviction risks discarding critical past tokens, leading to a slight performance decline. Thus, the choice of $W$ presents a trade-off between retaining local and global information.
Table 3: Performance of different setting of the window size $W$ on DeepSeek-R1-Distill-Llama-8B.
Dataset GSM8K (Budget=192)
FullKV W=4 W=8 W=16 W=25 W=32
Accurracy 74.68 68.69 70.13 70.28 73.61 73.18
Dataset MATH-500 (Budget=256)
FullKV W=8 W=16 W=32 W=52 W=64
Accurracy 62.0 54.4 54.6 55.2 57.6 57.2
The Selection of $\alpha$
We further explore the impact of the threshold $\alpha$ on the GSM8K dataset using two models. The value of $\alpha$ influences the MRI of each token, with $\alpha=0.0005$ producing optimal results for DeepSeek-R1-Distill-Llama-8B and $\alpha=0.0001$ yielding the best performance for DeepSeek-R1-Distill-Qwen-7B, as detailed in Table 4. First, when $\alpha$ is small, tokens are more readily flagged as important, but the MRI of these tokens is reduced, which does not accurately reflect the periodicity of their significance. Second, when $\alpha$ is large, the steps where tokens gain importance are not captured, preventing the detection of their recurring patterns.
Table 4: Performance of different setting of $\alpha$.
Model DS-Distill-Llama-8B (Budget=192)
FullKV $\alpha$=0.0001 $\alpha$=0.0005 $\alpha$=0.001
Accurracy 74.68 72.48 73.61 73.23
Model DS-Distill-Qwen-7B (Budget=768)
FullKV $\alpha$=0.00001 $\alpha$=0.0001 $\alpha$=0.001
Accurracy 91.74 90.44 90.75 90.67
Table 5: Performance of using different importance score.
LazyEviction 73.62
w/o the H1-Score 69.67
w/o the H2-Score 73.23
@@
We further investigate the influence of the MRI-Centric Importance Score, as shown in Table 5. If the importance score $\mathbf{I}_{t}$ calculation excludes the H1-Score, accuracy significantly decreases, underscoring the effectiveness of the recurrence interval tracking mechanism. Similarly, omitting the H2-Score leads to a slight accuracy drop, indicating that tokens with smaller MRI values are indeed more likely to become important in the future. Collectively, these findings demonstrate the efficiency of LazyEviction’s MRI-Centric Eviction Policy.
6 Limitations
While LazyEviction significantly improves KV cache efficiency for long reasoning tasks, several limitations remain. Firstly, our evaluation primarily focuses on mathematical reasoning, futher evaluation is required on programming and other types of reasoning datasets. Second, fixed-length observation windows limits LazyEviction’s flexibility across tasks and models. Future work should investigate adaptive window size strategies suitable for diverse reasoning scenarios. Third, although empirically validated, the conditions under which Token Importance Recurrence occurs lack formal theoretical characterization, particularly in complex reasoning pathways.
@@ 7 Conclusion
In this work, we address the critical bottleneck of KV cache memory overhead in long reasoning tasks. Our analysis reveals the Token Importance Recurrence phenomenon, where certain tokens receive renewed attention after multiple decoding steps. Our experiment shows the significant of these recurring tokens in reasoning tasks. Based on this insight, we propose LazyEviction, a novel lagged eviction framework that introduces temporal observation windows, recurrence interval tracking, and an MRI-centric eviction policy to preserve recurring tokens while dynamically managing memory usage. Experimental results demonstrate that LazyEviction achieves a 50% reduction in KV cache size while maintaining comparable performance, outperforming SOTA methods. By shifting from greedy eviction to predictive retention, our work provides a new paradigm for efficient KV cache compression in long-context generation scenarios.
Appendix A Technical Appendices and Supplementary Material
A.1 Mathematical Formulation of LazyEviction
@@
where $\mathbf{K}=[\mathbf{K}_{1},...,\mathbf{K}_{N_{t}}]\in\mathbb{R}^{N_{t}\times d_{h}}$, $\mathbf{V}=[\mathbf{V}_{1},...,\mathbf{V}_{N_{t}}]\in\mathbb{R}^{N_{t}\times d_{h}}$ and $d_{h}$ is the head dimension. We define $\mathbf{S}_{t}$ as the set of the KV pairs $(\mathbf{K}_{i},\mathbf{V}_{i})$ for all the previous tokens $i<N_{t}$ cached in GPU memory at decoding step $t$. Since the size of $\mathbf{K}$ and $\mathbf{V}$ grows linealy with the decoding step $t$, current methods try to evict the unimportant KV pairs from $\mathcal{S}_{t}$ at each step $t$. Unlike the previous KV compression methods, we make KV eviction decisions at decoding step $t=kW$ ($k\in\mathbf{N}^{+}$). Denote $\mathcal{S}^{*}_{t}$ as the optimal KV pairs selected from $\mathcal{S}_{t}$ under the cache budget $B$. Thus, the KV eviction problem can be formulate as:
${\mathcal{S}_{t}^{*}=\mathop{\arg\min}\limits_{\left|{\mathcal{S}_{t}^{\prime}}\right|=B,{\mathcal{S}_{t}^{\prime}}\subseteq{\mathcal{S}_{t}}}}&{{{\left\|{{\mathbf{A}_{t}}({\mathbf{Q}_{{N_{t}}}},\mathbf{K},\mathbf{V})-{\mathbf{A}_{t}^{\prime}}({\mathbf{Q}_{{N_{t}}}},\mathbf{K^{\prime}},\mathbf{V^{\prime}},{\mathcal{S}^{\prime}_{t}})}\right\|}_{2}}}\end{array}$ (4)
@@
It is intractable to directly solve Eq. 4 because of its combinatorial nature. We design a heuristic method to efficiently find the suboptimal solution for $\mathcal{S}^{*}_{t}$. First, as emphasized in previous works xiao2023efficient ; zhang2023h2o ; ghadia2025dialogue , retaining recent tokens is critical to ensuring coherence in the generated context. And to observe the importance variations of the newly generated tokens, we consistently preserve the latest $W$ KV pairs. Second, among the remaining past KVs, we retain $B-W$ KV pairs by prioritizing tokens most likely to exhibit importance recurrence in future steps. This selection is guided by an auxiliary score vector $\mathbf{I}_{t}$, which quantifies the potential future importance of tokens. Formally, our eviction policy is defined as:
@@
A.2 Experiment Setup
Models and Datasets
By fine-tuning with CoT data generated by DeepSeek-R1, even smaller distilled models can acquire long-reasoning capabilities similar to DeepSeek-R1. Thus, we select DeepSeek-R1-Distill-Llama-8B  [^3] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B and DeepSeek-R1-Distill-Qwen-7B  [^4] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B as the base models for evaluating KV cache compression. The evaluations are carried on two widely used mathematical reasoning benchmarks: GSM8Kcobbe2021training and MATH-500hendrycks2021measuring . GSM8K is a high-quality, linguistically diverse dataset of elementary-level math problems requiring multi-step reasoning and basic arithmetic operations to solve. MATH-500 contains 500 challenging problems drawn from high school math competitions, spanning domains such as algebra, geometry, and number theory.
Baselines
To validate the performance of LazyEviction, we compare it against Full Cache and three representative KV cache compression methods, including TOVA oren2024transformers , which retains tokens with the highest attention scores at each decoding step; H2O zhang2023h2o , which preserves heavy hitter tokens (identified by cumulative attention scores) and recent tokens during decoding; and RaaS hu2025efficient , which selects tokens with the newest timestamps at every decoding step.
Implementation Details
We implement LazyEviction in HuggingFace transformers library  wolf2020transformers to integrate it into the existing attention mechanism. We set the observation window size equal to the value of 80% MRI threshold based on the preliminary observations (Fig. 3(c)), and the number of recent tokens in H2O is equal to LazyEviction’s window size. Unless stated otherwise, we set $\alpha=0.0005$ for DeepSeek-R1-Distill-Llama-8B model and $\alpha=0.0001$ for DeepSeek-R1-Distill-Qwen-7B model. Experiments are conducted on NVIDIA V100 (32GB) GPUs. We use Ubuntu 20.04 with Linux kernel 6.8.0 and CUDA 12.8. For the budget setting in Table. 1, we first statistic the average length of different models on two datasets (shown in Table. 6). Then we adopt the compression ratio to compute the cache budget $B$.
Table 6: Token Length Statistics of Different Models
DeepSeek-R1-Distill-Llama-8B
GSM8K MATH-500
Prompt Output Length Total Length Prompt Output Length Total Length
85.97 257.79 343.76 95.82 583.7 679.52
DeepSeek-R1-Distill-Qwen-7B
Prompt Output Length Total Length Prompt Output Length Total Length
118.31 922.43 1040.74 128.26 2855.45 2983.71
A.3 Overall Procedure of LazyEviction
We provide the pseudo-code of LazyEviction in Algorithm. 1, which contains Recurrence Interval Tracking at each step $t$ and makeMRI-Centric Eviction when reach the decoding step $t=kW$ ($k\in\mathbf{N}^{+}$). Thus, Lazyeviction tasks the current decoding step $t$, observation window size $W$, cache budget $B$, $\alpha$, and $\mathbf{A}_{t}$, $\mathbf{K}$, $\mathbf{V}$ as input.

### v3

@@ 5 Experimental Results
In this section, we adopt variants of the DeepSeek-R1 distilled model [4]: DeepSeek-R1-Distill-Llama-8B (DS-Llama-8B) and DeepSeek-R1-Distill-Qwen-7B (DS-Qwen-7B), and Qwen-series reasoning model: Qwen3-4B [41] and QwQ-32B [42]. The evaluations are carried on five reasoning benchmarks from three different domains, including mathematical reasoning (GSM8K [5], MATH-500 [6], AIME [43]), Science QA (GPQA Diamond  [40]) and programming (LiveCodeBench [39]).
We compare LazyEviction with FullKV and four representative KV cache compression methods, including TOVA [13], H2O [16], RaaS [19], and R-KV [37]. We use the KV compression ratio $r$ to denote the KV cache budget across different baselines, where $r$ represents the proportion of KV cache usage relative to FullKV. More implementation details can be found in Appendix C.
@@ 5.1 Performance Comparison with Baselines
Table 1 and Table 2 present a comprehensive performance comparison of our LazyEviction and baselines. The results demonstrate that LazyEviction consistently outperforms other methods across all datasets, achieving performance comparable to the baseline FullKV with minimal deviation. In mathematical reasoning tasks, LazyEviction achieves performance close to FullKV with only 30%~50% KV budgets. Additionally, LazyEviction also achieves the best performance across GPQA Diamond and LiveCodeBench datasets, demonstrating its applicability among different domain tasks. It is worth noting that even though R-KV achieves performance close to LazyEviction in mathematical reasoning tasks, its performance significantly declines on other datasets. This is due to R-KV’s excessive reliance on the assumption that there are many similar tokens present in the reasoning path. In other tasks, similar tokens are not as prevalent as in mathematical tasks.
Table 1: Performance of our LazyEviction compared with baselines on mathematical reasoning dataset with DS-Llama-8B, DS-Qwen-7B, Qwen3-4B and QwQ-32B. The best results among all methods are in bolded.
Methods DS-Llama DS-Qwen Qwen3 QwQ
GSM8K (Compression Ratio $r=50\%$)
FullKV 81.73 89.92 93.32 95.61
RaaS 78.01 85.37 90.22 84.31
H2O 77.16 87.04 88.47 89.23
TOVA 72.25 80.52 85.13 72.10
R-KV 78.69 88.33 91.65 93.63
Ours 80.06 88.40 91.50 93.48
MATH-500 (Compression Ratio $r=50\%$)
FullKV 74.8 86.0 87.2 87.2
RaaS 71.2 82.4 84.0 80.4
H2O 67.2 80.4 79.6 80.4
TOVA 69.6 74.8 77.6 70.4
R-KV 73.2 83.6 83.4 85.2
Ours 75.2 85.4 85.8 85.4
AIME (Compression Ratio $r=30\%$)
FullKV 30.0 46.7 60.0 73.3
RaaS 23.3 36.7 46.7 53.3
H2O 26.7 33.3 40.0 53.3
TOVA 23.3 36.7 40.0 36.7
R-KV 26.7 43.3 46.7 56.7
Ours 30.0 43.3 53.3 66.7
Table 2: Performance of our LazyEviction compared with baselines on GPQA Diamond and LiveCodeBench with DS-Llama-8B and DS-Qwen-7B. The best results among all methods are in bolded.
Methods DS-Llama-8B DS-Qwen-8B
GPQA Diamond ($r=50\%$)
FullKV 37.4 55.7
RaaS 29.9 45.4
H2O 26.8 42.1
TOVA 30.9 44.3
R-KV 28.2 41.2
LazyEviction 36.9 54.6
LiveCodeBench ($r=40\%$)
FullKV 58.62 55.17
RaaS 53.45 39.66
H2O 53.45 50.00
TOVA 51.72 29.31
R-KV 48.28 46.55
LazyEviction 56.90 51.72
Trade-off between Accuracy and KV Cache.
We evaluate the performance of LazyEviction in optimizing the trade-off between KV cache budget and accuracy compared to baselines. As shown in Fig. 5, our method maintains higher accuracy across various budgets for both datasets and models. When the cache budget is large, other methods perform slightly worse than our approach. However, with a smaller cache budget, other methods experience substantial accuracy degradation. By implementing lagged KV eviction and leveraging MRI to track recurring tokens, our approach consistently achieves optimal performance, even under stringent budget constraints. Notably, on MATH-500 and AIME datasets, LazyEviction even outperformed FullKV’s performance while reducing the KV cache budget by 50%.
Figure 5: Trade-off between accuracy and KV Cache among different datasets and models.
@@ 5.2 Memory Efficiency of LazyEviction
Figure 6: KV cache memory usage of different algorithms with varying output length (0-8k tokens).
We evaluate the KV cache memory usage of different algorithms on DS-Qwen-7B model, analyzing its memory usage variation with output token length. As shown in Fig. 6, the memory consumption of FullKV increases linearly with the number of tokens. In contrast, the memory usage of TOVA, H2O, and RaaS algorithms remains constant once the token count exceeds the budget. LazyEviction, with the observation window mechanism, exhibits minor fluctuations in KV cache memory usage after surpassing the budget, but Fig. 6 indicates that its memory growth is minimal compared to the other three algorithms.
@@ 5.3 Ablation Study
Table 3: Baseline algorithms with observation window mechanism, evaluating on the GSM8K dataset using DS-Llama-8B with $r=50\%$ and $W=25$.
@@
LazyEviction 80.06
H2O 77.16
+ window 78.88 (+1.72)
TOVA 72.25
+ window 76.09 (+3.84)
RaaS 78.01
+ window 78.92 (+0.91)
@@
Since our method retains a slightly increased number of KV pairs within the observation window, which may lead to the perceived performance gains being predominantly attributed to these additional tokens. To eliminate the impact of the additional tokens within observation window, we explore the other algorithms with the observation window mechanism on performance. As shown in Table 3, when H2O, TOVA, and RaaS algorithms perform KV eviction every $W$ decoding steps, their performance improves. This improvement arises from the linear increase in retained KV pairs between consecutive decision decoding steps. However, their performance remains inferior to that of LazyEviction. LazyEviction employs MRI to track recurring tokens, effectively capturing their potential future importance, which leads to superior performance.
@@
We further investigate the influence of MRI-Centric Importance Score, as shown in Table 4. If the importance score $\mathbf{I}_{t}$ calculation excludes the H1-Score, accuracy significantly decreases, underscoring the effectiveness of the recurrence interval tracking mechanism. Similarly, omitting the H2-Score leads to a slight accuracy drop, indicating that tokens with smaller MRI values are indeed more likely to become important in the future. Collectively, these findings demonstrate the effectiveness of LazyEviction’s MRI-Centric Eviction Policy.
Table 4: Performance with different importance scores.
Model DS-Llama-8B DS-Qwen-7B
LazyEviction 80.06 88.40
w/o the H1-Score 76.11 (-3.95) 82.78 (-5.62)
w/o the H2-Score 79.67 (-0.39) 87.21 (-1.19)
@@ 6 Conclusion
In this work, we address the critical bottleneck of KV cache memory overhead in long reasoning tasks. Our analysis reveals the Token Importance Recurrence phenomenon, where certain significant tokens receive renewed attention after multiple decoding steps. Based on this insight, we propose LazyEviction, a novel lagged eviction framework that introduces temporal observation windows, recurrence interval tracking, and an MRI-centric eviction policy to preserve recurring tokens while dynamically reducing memory usage. Experimental results show that LazyEviction achieves performance close to or outperforms FullKV with only 30%~50% KV budgets, highlighting the importance of preserving recurring tokens on maintaining knowledge continuity in reasoning tasks. By shifting from greedy eviction to predictive retention, our work provides a new paradigm for efficient KV cache compression in long reasoning scenarios.
Limitations
Dynamically Adjusting in Observation Window Size.
We discuss in Appendix F.1 that the suitable observation window size is crucial for identifying recurring tokens. In Sec. 3, we find that the MRI distribution of recurring tokens varies across different models and tasks. Therefore, we propose to offline select 1% of samples to pre-statistically analyze the MRI distribution to determine the observation window size. However, this approach is inefficient and sub-optimal in practice. Thus, dynamically and adaptively adjusting the observation window based on task characteristics will be an important direction for improvement in LazyEviction in the future.
Evaluation Beyond Reasoning Tasks.
Although we have conducted experiments on tasks from different domains, our evaluation is still limited to reasoning tasks. This is because the core finding of this paper, the Token Importance Recurrence, is particularly significant in reasoning scenarios. We attempt to investigate TIR in non-reasoning scenarios. For example, we use Llama-3.1-8B-Instruct [^1] https:https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct to analyze the MRI distribution on the language modeling dataset C4222https://huggingface.co/datasets/allenai/c4. The results show that although TIR also exists, recurring tokens have a relatively smaller MRI (<10). In this case, the performance of LazyEviction is not significantly different from that of Cumulative Attention-based Eviction methods (e.g., H2O, RaaS). Therefore, we clarify that the observations and methods presented in this paper are more applicable to reasoning scenarios.
Evaluation on Larger Scale Models.
Our experiments cover reasoning models ranging from 4B to 32B in size. However, due to limitations in the experimental environment, we have not explored larger scale models, such as Qwen3-Max and DeepSeek-R1. Inference on models at the 100B level is extremely time- and resource-consuming. For example, evaluating 500 samples MATH-500 with DeepSeek-R1 requires days on 8 A100 GPUs. Nonetheless, we believe that our core finding, the Token Importance Recurrence phenomenon, and the importance of preserving recurring tokens exist across all reasoning models, and we will extend our experiments to more models in the future.
Appendix A Mathematical Formulation of LazyEviction
@@
where $\mathbf{K}=[\mathbf{K}_{1},...,\mathbf{K}_{N_{t}}]\in\mathbb{R}^{N_{t}\times d_{h}}$, $\mathbf{V}=[\mathbf{V}_{1},...,\mathbf{V}_{N_{t}}]\in\mathbb{R}^{N_{t}\times d_{h}}$ and $d_{h}$ is the head dimension. We define $\mathcal{S}_{t}$ as the set of the KV pairs $(\mathbf{K}_{i},\mathbf{V}_{i})$ for all the previous tokens $i<N_{t}$ cached in GPU memory at decoding step $t$. Since the size of $\mathbf{K}$ and $\mathbf{V}$ grows linealy with the decoding step $t$, current methods try to evict the unimportant KV pairs from $\mathcal{S}_{t}$ at each step $t$. Unlike the previous KV compression methods, we make KV eviction decisions at decoding step $t=kW$ ($k\in\mathbf{N}^{+}$). Denote $\mathcal{S}^{*}_{t}$ as the optimal KV pairs selected from $\mathcal{S}_{t}$ under the cache budget $B$. Thus, the KV eviction problem can be formulate as:
$\mathcal{S}_{t}^{*}=\mathop{\arg\min}\limits_{\left|{\mathcal{S}_{t}^{\prime}}\right|=B,{\mathcal{S}_{t}^{\prime}}\subseteq{\mathcal{S}_{t}}}\\ {\left\|{\mathbf{A}_{t}}({\mathbf{Q}_{{N_{t}}}},\mathbf{K},\mathbf{V})-{\mathbf{A}_{t}^{\prime}}({\mathbf{Q}_{{N_{t}}}},\mathbf{K^{\prime}},\mathbf{V^{\prime}},{\mathcal{S}^{\prime}_{t}})\right\|}_{2}$ (4)
@@
It is intractable to directly solve Eq. 4 because of its combinatorial nature. We design a heuristic method to efficiently find the suboptimal solution for $\mathcal{S}^{*}_{t}$. First, as emphasized in previous works [12; 16; 20], retaining recent tokens is critical to ensuring coherence in the generated context. And to observe the importance variations of the newly generated tokens, we consistently preserve the latest $W$ KV pairs. Second, among the remaining past KVs, we retain BWB-W KV pairs by prioritizing tokens most likely to exhibit importance recurrence in future steps. This selection is guided by an auxiliary score vector $\mathbf{I}_{t}$, which quantifies the potential future importance of tokens. Formally, our eviction policy is defined as: